Calculating Weighted Average Cost of Capital and Economic Value Added (EVA)

Ignacio, Inc., had after-tax operating income last year of $1,195,000. Three sources of financing were used by the company: $2 million of mortgage bonds paying 4 percent interest, $5 million of unsecured bonds paying 6 percent interest, and $10 million in common stock, which was considered to be relatively risky (with a risk premium of 8 percent). The rate on long-term treasuries is 3 percent. Ignacio, Inc., pays a marginal tax rate of 30 percent.

Required:

1. Calculate the after-tax cost of each method of financing. Enter your answers as decimal values rounded to three places. For example, 4.36% would be entered as ".044".

Mortgage bonds
Unsecured bonds
Common stock

2. Calculate the weighted average cost of capital for Ignacio, Inc. Round intermediate calculations to four decimal places. Round your final answer to four decimal places before converting to a percentage. For example, .06349 would be rounded to .0635 and entered as "6.35" percent.

%

Calculate the total dollar amount of capital employed for Ignacio, Inc.

$

3. Calculate economic value added (EVA) for Ignacio, Inc., for last year. If the EVA is negative, enter your answer as a negative amount.

$

Is the company creating or destroying wealth?

Destroying

4. What if Ignacio, Inc., had common stock which was less risky than other stocks and commanded a risk premium of 5 percent? How would that affect the weighted average cost of capital?
Lower

What is the new EVA? In your calculations, round weighted average percentage cost of capital to four decimal places. If the EVA is negative, enter your answer as a negative amount.

$
Answer #1

1) Calculation of payout of each method of Finance:

Rate of dividend on common stock = risk free rate + risk premium

= 0.03+0.08 = 0.11 i.e. 11%

Particulars Amount (Financed) Dividend / Interest After Tax Cost in amount After tax Cost of capital
(a) (b) (c) (d) (e)
Mortgage Bond 2,000,000 80,000 56,000 0.0280
Unsecured Bond 5,000,000 300,000 210,000 0.0420
Common Stock 10,000,000 1,100,000 1,100,000 0.11

2) Calculation of weighted average cost of capital:

Particulars Amount (Financed) Weight After Tax Cost in amount After tax Cost of capital Weighted average cost (WACC)
(a) (b) (c) (d) (e) (f = e*c)
Mortgage Bond 2,000,000 2/17 56,000 0.0280 0.00329
Unsecured Bond 5,000,000 5/17 210,000 0.0420 0.01235
Common Stock 10,000,000 10/17 1,100,000 0.11 0.06470
Total 17,000,000 0.08034

Therefore, weighted average cost of capital is 0.08034 i.e. 8.034%

3) Economic Value Added:

EVA = NOPAT - (WACC* Capital employed)

EVA = 1,195,000 - (0.08034*17,000,000)

EVA = (170780)

As the EVA figure is negative, company is destroying the wealth of the investors instead of creating it.

4) If the company's common stocks were less risky and had commanded only 5% risk premium, then,

Rate of dividend on common stock = risk free rate + risk premium

= 0.03+0.05 = 0.08 i.e. 8%

Calculation of weighted average cost of capital:

Particulars Amount (Financed) Weight After Tax Cost in amount After tax Cost of capital Weighted average cost (WACC)
(a) (b) (c) (d) (e) (f = e*c)
Mortgage Bond 2,000,000 2/17 56,000 0.0280 0.00329
Unsecured Bond 5,000,000 5/17 210,000 0.0420 0.01235
Common Stock 10,000,000 10/17 800,000 0.08 0.04706
Total 17,000,000 0.0627

Therefore the company's WACC would have been = 0.0627 i.e. 6.27%

EVA would have been = 1,195,000 - (0.0627*17,000,000)

= 129100

Had the company's stock been less risky, company would have created a positive economic value added for its investors.